The length of a rectangle is 7 inches less than 3 times the width. If the area of the rectangle is 40 square inches, find the dimensions.
Answer by ewatrrr(24785)


Hi
The length of a rectangle is 7 inches less than 3 times the width
Let x and (3x-7)represent the width and length respectively
Question states***
x(3x-7) = 40in^2
3x^2 -7x - 40 = 0
factoring
(3x +8)(x-5)=0 Note:SUM of the outside terms(-15x) and the inside terms(8x) = -7x
(3x +8)=0 |tossing out negative solution for length
(x-5)=0 x = 5", the width. the length is 8" (3*5"-7")